OpenStudy (anonymous):

In triangle ABC , A=32 degrees, c=8 and b=7. find C A. 110.7° B. 20.7° C. 30.4° D. 120.4°

OpenStudy (anonymous):

solve for side a using the cosine law: \(\large a^2=b^2+c^2-2\cdot b\cdot c\cdot cosA \) after getting side a, use that in the sine law to solve for angle C: \(\large \frac{sinA}{a}=\frac{sinC}{c} \)